OpenStudy (jaweria):

calculate the volume of 10.0 mol L-1 nitric acid solution required to prepare 500.0 mL of 0.500 mol L-1 solution

OpenStudy (danjs):

been a while since chem, but isnt the dilution someting like concentration 1 * Volume 1 = Concentration 2 * Volume 2

OpenStudy (sweetburger):

\[M_1V_1=M_2V_2\]

OpenStudy (sweetburger):

Make sure you have concentration in units of molarity (mol/L) and volume in units of Liters.

OpenStudy (jaweria):

could you guys solve this now?

OpenStudy (danjs):

you use the formula from above, keeping units consistent, using liters concentration 1 * Volume 1 = Concentration 2 * Volume 2 10.0 mol/L * (Volume 1) = 0.500 mol/L * 0.500 L 10.0 * V = 0.500 * 0.500